OpenStudy (anonymous):

Verify Rolle’s theorem for the function f where f(x) = sin x, x Î [-2π, 2π].

Parth (parthkohli):

Continuous in \([-2\pi, 2\pi]\)? Yes. Differentiable in \((-2\pi, 2\pi)\)? Yes. \(f(2\pi) = f(-2\pi)\)? Indeed.

Parth (parthkohli):

So when exactly is the slope of the tangent zero?\[f'(x) = \cos (x) = 0 \Rightarrow x=\pm \pi/4, \pm 3\pi /4 \]

OpenStudy (anonymous):

OK. Anything else Sir?

Parth (parthkohli):

That's very much it.

OpenStudy (anonymous):

OK. Thanks

OpenStudy (anonymous):

f'(x)=cosx=0 x=+-pi/4

OpenStudy (anonymous):

,+-3pi/4

OpenStudy (anonymous):

please can you explain that line for me ?

OpenStudy (anonymous):

@perl

OpenStudy (perl):

cos x = 0 when x = pi/2 , 3pi/2 , 5pi/2, etc

OpenStudy (anonymous):

thanks bosssssssssssssssssssssssssssssssssss

OpenStudy (perl):

I think parth made a typo above

OpenStudy (anonymous):

yes. thanks for the correction

OpenStudy (anonymous):

yap. i got that.

OpenStudy (perl):

Continuous in \( [-2\pi, 2\pi] \) ? Yes. Differentiable in \((-2\pi, 2\pi)\)? Yes. \(f(2\pi) = f(-2\pi)\)? Indeed. So when exactly is the slope of the tangent zero?\[f'(x) = \cos (x) = 0 \Rightarrow x=\pm \pi/\color{red}{2}, \pm 3\pi /\color{red}{2}\]

OpenStudy (anonymous):

perl, is there a shorter way to prove continuity and derivability ?

OpenStudy (perl):

well, sin x and cos x are continuous and differentiable for all real numbers

OpenStudy (perl):

you don't need to prove that here. we can assume that it has been proved earlier , and it is a theorem

OpenStudy (perl):

Continuous in \( [-2\pi, 2\pi] \) ? Yes because it is continuous for all reals. Differentiable in \((-2\pi, 2\pi)\)? Yes. because it is differentiable for all reals. \(f(2\pi) = f(-2\pi)\)? Indeed. So when exactly is the slope of the tangent zero?\[f'(x) = \cos (x) = 0 \Rightarrow x=\pm \pi/\color{red}{2}, \pm 3\pi /\color{red}{2}\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

can you prove this if continous and derivable ? \[\left| x \right| +\left| x+1 \right| for x=0 and x=1\]

OpenStudy (perl):

can you prove this if continuous and derivable ? \[ \Large {\left| x \right| +\left| x+1 \right| ~for~ x=0,1}\]

OpenStudy (anonymous):

yes

OpenStudy (perl):

\( \large f(x) = |x | \) is continuous and \( \large g(x) = |x+1 | \) is continuous, and the sum of two continuous functions is continuous, so \( \large f(x) + g(x) \) is continuous .

OpenStudy (anonymous):

ok

OpenStudy (perl):

now for the differentiable part, lets define the absolute value function in terms of cases , or piecewise define it

OpenStudy (anonymous):

waw.

OpenStudy (perl):

$$ \Large{ |x| + |x+1|= \begin{cases} x + (x+1) & \text{if } x \geq 0\\ -x + (x+1) & \text{if } -1\leq x<0\\ -x - (x+1)& \text{if } x<-1 \end{cases}} $$ simplified this gives us $$\Large{ |x| + |x+1|= \begin{cases} 2x+1 & \text{if } x \geq 0\\ 1 & \text{if } -1\leq x<0\\ -2x - 1& \text{if } x<-1 \end{cases} } $$

OpenStudy (anonymous):

pk

OpenStudy (anonymous):

ok got that

OpenStudy (perl):

$$ \Large{ |x| + |x+1|= \begin{cases} 2x+1 & \text{if } x \geq 0\\ 1 & \text{if } -1\leq x<0\\ -2x - 1& \text{if } x<-1 \end{cases} } \\{} \\ \Large{ (|x| + |x+1|) '= \begin{cases} 2 & \text{if } x > 0\\ undefined & \text{if } x = 0\\ 0 & \text{if } -1< x<0\\ undefined & \text{if } x = -1\\ -2 & \text{if } x<-1 \end{cases} } $$

OpenStudy (perl):

do you see why the derivative is undefined at x = 0 and x = -1, because the slope changes depending on which direction you are coming from, from the right or left of x = 0 , right or left of x = -1. The slopes do not match at the point .

OpenStudy (anonymous):

waw. this is great

OpenStudy (anonymous):

are you a lecturer ?

OpenStudy (perl):